OpenStudy (anonymous):

A house contains 46 bricks along its bottom row. If each row above decreases by 4 bricks, how many bricks are on the 12th row?

OpenStudy (amistre64):

A1 = 46 A2 = 46 - 4(1) A3 = 46 - 4(2) An = 46 - 4(n-1) perhaps?

OpenStudy (amistre64):

A12 = 46 - 4(11) = 2

OpenStudy (anonymous):

Thank you! This is the same answer I got - just double checking which forula to use - so I must be correct :). Thanks again!

OpenStudy (amistre64):

youre welcome; i got no idea what formula to use, just tried a little of me uncommon sense ;)

OpenStudy (anonymous):

I think it is arithmetic series formula. I'll double check. Bye and thanks!